OpenStudy (anonymous):

Find real and imaginary parts of (a+3bi)^4

OpenStudy (anonymous):

an imaginary number could change to a radius with an angle. change it then power the radius by 4 and multiply the angle with four . i wish it would help you

OpenStudy (anonymous):

i just need to solve this

OpenStudy (anonymous):

plz expand this formula i can do the rest

OpenStudy (anonymous):

r= sqrt(x^2+y^2) phi=tan -1 ( y/x) x = real part y= imaginary part

OpenStudy (anonymous):

this just needs to be solved!!! the R & I parts will show up

OpenStudy (anonymous):

????????????????????????????????????????????????

OpenStudy (anonymous):

so what ?

OpenStudy (anonymous):

i've solved it using binomial theorem by myself!!

OpenStudy (anonymous):

thanks to

OpenStudy (anonymous):

myself!! :-/

OpenStudy (anonymous):

:D